### v3

@@ Abstract
This paper develops a dynamic optimization framework for evaluating the optimal longevity of a dynasty under a critical-level utilitarian social welfare criterion. Unlike standard models that assume an infinite planning horizon and apply discounting to future generations, we endogenize the planning horizon itself and examine the welfare-maximizing number of generations in a roundabout production economy with capital accumulation. The model incorporates a well-being threshold that must be met for generational existence to be considered ethically justified, providing an alternative to traditional utilitarian aggregation. Using finite-horizon dynamic programming, we derive closed-form consumption trajectories and analyze how technological parameters and intertemporal preferences affect the optimal generational span. Our results show that the optimal horizon need not be infinite—even under zero discounting—and that intergenerational equity may require limiting dynasty length.
@@
O21 I30 C61
Keywords:
Critical-level Utilitarianism Dynamic Programming Optimal Population Intergenerational Equity
@@ 1 Introduction
Economic models of intertemporal planning often assume an infinite horizon, implicitly treating dynasties or economies as perpetually forward-looking entities. However, such assumptions can obscure important trade-offs between current and future welfare, particularly when the viability of future generations depends on limited resources and capital dynamics. In this paper, we examine a social planner’s problem where the planning horizon—the number of generations considered—is endogenously chosen. By adopting a critical-level utilitarian welfare function, we evaluate the contribution of each generation relative to a minimum well-being threshold. This framework allows us to study how productivity, capital accumulation, and time preference interact to determine the ethically and economically optimal duration of a dynasty.
We begin by positing a (classical) unitilitarian social planner whose objective is to indentify a state $x$ that maximizes the sum of the utilities $\mathcal{V}=\sum_{i\in n(x)}\Upsilon_{i}(x)$, which we refer to as the population value, where $n(x)$ denotes the set of individuals and $\Upsilon_{i}(x)$ denotes the lifetime utility (i.e., the level of well-being) of an individual $i$. For simplicity, let us assume that all individuals are identical and consume an equal share, denoted by $c$, from a fixed amount of resource $R>0$. A utilitarian axiology examines the problem of finding the population size $n=R/c$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\Upsilon(c_{i})=n\Upsilon(c=R/n)$. The utility function $\Upsilon(c)$ is assumed to be monotonically increasing and concave. In addition, with respect to the population value $\mathcal{V}=n\Upsilon(c)$, zero utility ($\Upsilon=0$) is equivalent to the absence of life ($n=0$).
The consumption level $c=\nu$, where $\Upsilon(\nu)=0$, be referred to as the well-being subsistence level, following Dasgupta (2019). Owing to the monotonicity of the utility function, $\Upsilon(c)>0$ must hold as long as $c>\nu$ (see Figure 1 (left)). It then follows that the domain allowed for $c$ is $(\nu,\infty)$ since $\Upsilon(c)$ represents the level of attainable lifetime utility. Figure 1 (left) depicts the optimal solution $c=\omega$, which satisfies the first-order condition for maximizing the population value, i.e., $\Upsilon^{\prime}(\omega)=\Upsilon(\omega)/\omega$. Owing to the concavity of the utility function, if the well-being subsistence approaches zero ($\nu\to 0$), as Córdoba (2023) refers to the homothetic case, the optimal consumption also approaches zero ($\omega\to\nu$), while the optimal size of the population approaches infinity ($n^{*}=R/\omega\to\infty$), leading to a conclusion that is considered repugnant (Parfit, 1986).
Meanwhile, Blackorby et al (2005) introduced a population axiology known as critical-level utilitarianism. Within this framework, a social planner aims to determine a set of consumption schedules $c_{i}$ that maximizes the population value $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)$, where an individual’s contribution to the population value is evaluated relative to the critical utility level $\alpha$ (see Figure 1 (right)). That is, an individual $i$’s life contributes positively to the population value if $c_{i}>\kappa=\Upsilon^{-1}(\alpha)$ but negatively otherwise. Note that setting $\alpha>0$ prevents the repugnant conclusion even when the underlying utility function is homothetic ($\nu=0$). For convenience, we employ a utility function of the form $\Upsilon(c_{i})=\log c_{i}+\alpha$ for analyses so that the value function for the critical-level utilitarian population value in our maximization study becomes $\mathcal{V}=\sum_{i=1}^{n}\left(\Upsilon(c_{i})-\alpha\right)=\sum_{i=1}^{n}\log c_{i}$.
As we recall our primary idea of incorporating roundabout production into the optimal population framework, let us introduce (discrete) time and assume that the number of individuals per period is fixed to one unit; that is, all (representative) individuals are identified by the period $t$ during which they live. [^1] We may therefore refer to $t$ as a generation. The population value function then becomes $\mathcal{\mathcal{V}}=\sum_{t=0}^{n}\log c_{t}$. Note that this type of value function was widely employed in the previous Ramsey–Cass–Koopmans type of dynastic optimization model (including Nordhaus, 1992; Stern, 2007), except that the future generations were discounted according to the temporal deviation from the present, and typically, the number of potential generations $n$ remains unquestioned (and set to infinity). In contrast, in this study, potential generations are accounted for equally, and the number of generations is optimized.
The remainder of the paper proceeds as follows. In the following section, we first describe the model of roundabout production on the basis of Cobb–Douglas technology. We then define our main problem in the form of finite horizon dynamic programming, where the objective function comprises a series of Bernoullian contributions, and analytically derive a general solution. In Section 3, we introduce two broad settings of parameters to simplify the general solution and allow us to study the optimal trajectory of the variables in a numerical fashion. The first setting is the AK setting, where the output elasticity of capital is set to unity. The second is the zero discounting (ZD) setting, where the discount factor is set to unity. We also study the peculiarity of the two models with respect to intergenerational equity and the optimal initial action of consumption. Section 4 concludes the paper.
@@ 2.1 Roundabout production
where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$. For the relevant parameters, $10<\theta<$ 1 denotes the output elasticity of capital. The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span. Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity. [^2] Otherwise, $B$ is referred to as the dual productivity. See Appendix 1 for more details.
@@
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively. Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$.
@@ 2.2 Finite horizon dynamic programming
The primary problem here is to solve for an optimal consumption path given a planning horizon. The Bellman equation of the problem is as follows:444 Note that the objective function of the primary problem (2.1) is given at $t=0$, i.e., $\mathcal{V}[n]=\mathcal{V}_{0}[k_{0};n]$.
@@
We hereafter aim to maximize this function with respect to the planning horizon $n$.
Our approach to analyzing the population value (8) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[n]$ with respect to $n$ does not seem to provide meaningful insights. In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[n]$ and the population value $\mathcal{V}[n]$, for any given planning horizon $n$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$. We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings.
Table 1 summarizes the parameter settings chosen for the numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$. Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
@@
case ${A}$ $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $n^{*}$ $\mathcal{V}[n^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ $1$ $+$ $\infty$ $84.7$
II $1.01$ $0.992$ $1$ $+$ $95$ $60.0$
III $1/\beta$ $0.992$ $1$ 00 $73$ $55.6$
IV $1.005$ $0.992$ $1$- $58$ $51.0$
V $1.05$ $1$ $0.992$ $+$ $\infty$ +$\infty$+ ∞
VI $1.2$ $1$1 00 $(281)$ $41.7$
VII $1.05$ $1$ $0.991$- $117$$\infty$- ∞
VIII $1$ $1$ $1$ 00 00 $54$ $55.2$$\infty$- ∞
IX $1/\beta$ $0.992$2 00 00 $53$ $49.1$ $-1843.2$
@@
    $\theta\approx 0.955392$ where 1.2($1{1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=$ 1.
@@ 3.1 AK setting
The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $n\to\infty$ is an optimal policy. The optimal consumption path (7) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For the sake of the analysis, let us take the derivative with respect to ${n}$.
@@
In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$. Therefore, assuming that ${A}>1$ is relevant in this setting.555 Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
Figure 2 (left) depicts the functions $f[n]$ and $g[n]$ under different parameters for cases I–IV. Clearly, $f[n]$ is a linear function whose slope is $\log({A}\beta)$, and $g[n]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$ and differentiate the productivity ${A}$ at four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[n]$ will intersect with $g[n]$ at a single point, and the population value function $\mathcal{V}[n]$ has a single peak. If ${A}>1/\beta$ so that $f[n]$ has a positive slope, then $f[n]$ and $g[n]$ could either intersect with two points where the population value $\mathcal{V}[n]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $n$. Figure 2 (right) depicts the population value functions for cases I–IV.
@@
Figure 3: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right). The parameters corresponding to each of the cases are given in Table 1.
@@ 3.2 ZD setting
where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following zero discount verification of the optimal trajectory for undiscounted contributions:
@@
The above result indicates that if $B>1$, then the population value is ever increasing, and $n\to\infty$ must be the optimal solution. However, if $B<1$, then $n\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $n^{*}\ll\infty$. The case where $B=1$ is the knife-edge case in which the planning horizon $n$ does not matter (beyond a certain length) in maximizing the population value. Case V corresponds to $B>1$, and $n^{*}=\infty$; case VII corresponds to $B<1$, and $n^{*}\ll\infty$; and case VI corresponds to $B=1$, and $n^{*}$ is any number greater than $281$.777 The number is subject to decimal rounding. Figure 4 depicts the population value function $\mathcal{V}[n]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[n]$ and the capital intensity $k_{t}^{*}[n]$, given a planning horizon $n$, for sample ZD settings V, VI, and VII.
The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (13). Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $n^{*}\to\infty$. If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
@@
Figure 4: Population value function (top), trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). Note that the population value function of the top-left panel is increasing indefinitely. The parameters corresponding to each of the cases are given in Table 1.

### v4

@@ 1 Introduction
Standard economic models of optimal growth typically assume an infinite planning horizon. This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey1928). While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints. In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (DASGUPTA2019). By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (PARFIT).
This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level). An individual’s existence contributes to social welfare only if their utility exceeds this critical level; otherwise, it is considered a negative contribution. By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
Methodologically, our approach establishes a structural isomorphism between static population ethics and dynamic growth theory. In the tradition of optimal population theory initiated by Samuelson1975, a social planner typically chooses the optimal population size $N$ to maximize social welfare under resource constraints. In our dynamic framework, we reinterpret this problem by viewing the number of generations $N$ (longevity) as the choice variable. This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin. Consequently, our finite-horizon solution implies a conscious decision to cease reproduction, rather than a failure of time consistency. It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (AsheimZuber2014). RDU prevents the conclusion by strictly discounting the value of additional lives. However, this approach can lead to the opposite extreme—an “anti-population bias” or “reverse repugnant conclusion”—where a small population with high utility is preferred over a larger population that is also well-off. Furthermore, in a dynamic context, RDU implies that the welfare weight of future generations depends on their relative rank in history. This dependency destroys the recursive structure of the optimization problem and leads to time inconsistency. In contrast, CLU provides a transparent, absolute threshold for “lives worth living” that preserves both intergenerational equity (for those above the threshold) and the tractability of dynamic programming.
The remainder of the paper proceeds as follows. Section 2 describes the model of roundabout production using Cobb–Douglas technology and defines the main problem as finite horizon dynamic programming. We analytically derive the optimal consumption path and the optimal longevity $N^{*}$. In Section 3, we analyze the properties of the solution under two parameter settings: the AK setting (unity output elasticity of capital) and the Zero Discounting (ZD) setting (intergenerational equity). We show that under specific conditions—particularly when productivity is low—the ethically optimal policy is to choose a finite horizon, suggesting that sustainability does not always imply indefinite continuation. Section 4 concludes with discussions on the ethical implications of our findings for modern demographic trends.
2 Model
2.1 Critical-Level Preferences
Before describing the production technology, we first specify the social planner’s objective function. We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$. The planner aims to determine the consumption path and the planning horizon (number of generations) to maximize the following social welfare function:
$\mathcal{V}=\sum_{t=0}^{N}\left(\Upsilon(c_{t})-\alpha\right)$ (1)
where $\alpha$ represents the critical level of utility. An individual’s life contributes positively to social welfare if and only if their utility exceeds this critical threshold $\alpha$; otherwise, it reduces social welfare.
For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$. This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus1992; Stern2007), which facilitates comparisons with standard benchmarks. Under this specification, the contribution of generation $t$ to the population value becomes:
$\Upsilon(c_{t})-\alpha=(\log c_{t}+\alpha)-\alpha=\log c_{t}$ (2)
Consequently, the objective function simplifies to $\mathcal{V}=\sum_{t=0}^{N}\log c_{t}$. We define the well-being subsistence level $\nu$ as the consumption level where utility is exactly zero, i.e., $\Upsilon(\nu)=0$. In our logarithmic specification, this corresponds to $\nu=e^{-\alpha}$. As illustrated in Figure 1, the domain of permissible consumption is $c>\nu$. The critical level $\alpha$ thus serves as a “buffer” above the biological subsistence level. The planner will only choose to extend the dynasty if the consumption level allows for a utility strictly greater than the critical level (i.e., $\log c_{t}>0$, or $c_{t}>1$ in normalized terms), ensuring that each generation enjoys a life worth living.
@@ 2.2 Roundabout production
where $Y_{t}$ denotes the economy’s output, $K_{t}$ denotes capital, and $L_{t}$ denotes labor, all of which are effective during period $t$. For the relevant parameters, $0<\theta<1$ denotes the output elasticity of capital. The level of technology is indicated by productivity, denoted by $A$, which is assumed to be fixed throughout the considered time span. Note that ${B}\equiv{A}\theta^{\theta}(1-\theta)^{1-\theta}$ is the cost function-based productivity. [^1] Otherwise, $B$ is referred to as the dual productivity. See Appendix 1 for more details.
@@
where $k_{t}=K_{t}/L$ and $k_{t+1}=K_{t+1}/L$ denote capital intensities in periods $t$ and $t+1$, respectively. Additionally, $y_{t}=Y_{t}/L$ and $c_{t}=C_{t}/L$ denote per capita output and consumption, respectively, in period $t$. The social planner aims to maximize the critical-level utilitarian population value $\mathcal{\mathcal{V}}$ as previously specified, subject to the state transition function (5), i.e.,
@@
[…]
given the initial state $k_{0}$, we assume complete depreciation $\delta=1$ and introduce the discount factor $\beta\leq 1$. We let $\beta\to 1$ for a (critical-level) utilitarian assessment. [^2] The discount factor $\beta$ is otherwise referred to as the rate of time preference. In the context of dynastic optimization, $\beta$ may be referred to as the rate of generational preference (of the population). The optimal consumption path $(c_{0}^{*},c_{1}^{*},\cdots,c_{N}^{*})$ is clearly dependent on the planning horizon $N$. We therefore solve the above problem hierarchically. That is, we first solve for the optimal consumption path given $N$ via finite horizon dynamic programming and obtain the population value function with respect to the planning horizon $\mathcal{\mathcal{V}}[N]$; therefore, we search for the optimal planning horizon $N^{*}$.
@@ 2.3 Finite horizon dynamic programming
The primary problem here is to solve for an optimal consumption path given a planning horizon. The Bellman equation of the problem is as follows: [^3] Note that the objective function of the primary problem (6a) is given at $t=0$, i.e., $\mathcal{V}[N]=\mathcal{V}_{0}[k_{0};N]$.
@@
We hereafter aim to maximize this function with respect to the planning horizon $N$. Our approach to analyzing the population value (11) adopts a numerical rather than an analytical framework because the derivative of $\mathcal{V}[N]$ with respect to $N$ does not seem to provide meaningful insights. In the following section, the optimal trajectory of contributions $\log c_{t}^{*}[N]$ and the population value $\mathcal{V}[N]$, for any given planning horizon $N$, becomes manageable under $\theta=1$ (known as the AK setting) and $\beta<1$. We also find that the trajectory of contributions $\log c_{t}^{*}[\infty]$ and the population value $\mathcal{V}[\infty]$ for an infinite planning horizon is evaluable under $\beta\theta<1$. We therefore base our study of ZD ($\beta=1$) on this parameter setting (i.e., $\beta\theta<1$, which indicates that $\theta<1$, consistent with a Cobb–Douglas model). In the following section, we delve into the abovementioned two broad settings of parameters, namely, AK production with future discounting ($\theta=1$ and $\beta<1$), which we term AK settings, and Cobb–Douglas production without future discounting ($\theta<1$ and $\beta=1$), which we term ZD settings. Table 1 summarizes the parameter settings chosen for the numerical examinations. Note that cases I–IV correspond to AK settings whose solution paths are characterized by the sign of $\log(A\beta)$, whereas cases V–VII correspond to ZD settings whose solution paths are characterized by the sign of $\log B$. Cases VIII and IX correspond to the parameter settings where $\log({A}\beta)=\log B=0$.
@@
case A{A} $\beta$ $\theta$ $\log({A}\beta)$ $\log B$ $N^{*}$ $\mathcal{V}[N^{*}]$ $\mathcal{V}[\infty]$
I $1.012$ $0.992$ 11 ++ $\infty$ $84.7$
II $1.01$ $0.992$ 11 ++ $95$ $60.0$
III $1/\beta$ $0.992$ 11 0 $73$ $55.6$
IV $1.005$ $0.992$ 11 - $58$ $51.0$
V $1.05$ 11 $0.992$ ++ $\infty$ $\infty$
VI $1.2$ 11 *1 0 $(281)$ $41.7$
VII $1.05$ 11 $0.991$ - $117$ -$\infty$
VIII 11 11 11 0 0 $54$ $55.2$ -$\infty$
IX $1/\beta$ $0.992$ *2 0 0 $53$ $49.1$ 1843.2-1843.2
@@
    $\theta\approx 0.955392$ where ${1.2}(1-\theta)^{1-\theta}\theta^{\theta}=B=1$.
@@ 3.1 AK setting
The AK model, which was formally developed by Frankel, is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy. The optimal consumption path (10) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For the sake of the analysis, let us take the derivative with respect to N{N}.
@@
In AK models with complete capital depreciation, $\theta=\delta=1$ leads to ${A}=1+\rho$. Therefore, assuming that A>1{A}>1 is relevant in this setting. [^4] Note also that $A\beta=(\delta+\rho)\beta$ indicates the discrepancy between the gross interest rate and generational preference rate of the population.
@@
Figure 2 (left) depicts the functions $f[N]$ and $g[N]$ under different parameters for cases I–IV. Clearly, $f[N]$ is a linear function whose slope is $\log({A}\beta)$, and $g[N]$ monotonously increases and approaches zero, i.e., $g[\infty]=0$. Here, we fix the discount factor $\beta$ and differentiate the productivity A{A} at four different levels. As long as ${A}\leq 1/\beta$, so that $\log({A}\beta)\leq 0$, i.e., the slope is zero or negative, $f[N]$ will intersect with $g[N]$ at a single point, and the population value function $\mathcal{V}[N]$ has a single peak. If ${A}>1/\beta$ so that $f[N]$ has a positive slope, then $f[N]$ and $g[N]$ could either intersect with two points where the population value $\mathcal{V}[N]$ may rise, fall and then rise again, or never intersect so that the population value may rise indefinitely with respect to the planning horizon $N$. Figure 2 (right) depicts the population value functions for cases I–IV.
@@
Figure 3: Population value function (top), seleted trajectories of undiscounted contribution (middle) and capital intensity (bottom), for AK setting cases I (left), III (center), and IV (right). The trajectories are selected for $N=200,400,600$.
@@ 3.2 ZD setting
where the optimal trajectory for undiscounted contributions is geometrically convergent. We can then apply $\beta=1$ to arrive at the following:
@@
The above result indicates that if $B>1$, then the population value is ever increasing, and $N\to\infty$ must be the optimal solution. However, if $B<1$, then $N\to\infty$ must not be optimal, and the population value must be maximized at a finite horizon $N^{*}\ll\infty$. The case where $B=1$ is the knife-edge case in which the planning horizon $N$ does not matter (beyond a certain length) in maximizing the population value. Case V corresponds to $B>1$, and $N^{*}=\infty$; case VII corresponds to $B<1$, and $N^{*}\ll\infty$; and case VI corresponds to $B=1$, and $N^{*}$ is any number greater than 281281.666The number is subject to decimal rounding. Figure 4 depicts the population value function $\mathcal{V}[N]$, along with the optimum trajectory of undiscounted contribution $\log c_{t}^{*}[N]$ and the capital intensity $k_{t}^{*}[N]$, given a planning horizon $N$, for sample ZD settings V, VI, and VII. The final case VIII relates to the ZD/AK setting ($\theta=\beta=1$), where $B=A=1+\rho>1$, under complete depreciation, as described in (16). Thus, if $\rho>0$, then $A>1$, in which case the population value would increase without bound, as described above, leading to $N^{*}\to\infty$. If, in turn, $A=1+\rho=1$, as in case VIII, the infinite horizon population value is evaluated as follows:
@@
Figure 4: Population value function (top), selected trajectories of undiscounted contribution (middle) and capital intensity (bottom), for ZD setting cases V (left), VI (center), and VII (right). The trajectories are selected for $N=200,400,600$. Note that the population value function of the top-left panel is increasing indefinitely. The parameters corresponding to each of the cases are given in Table 1.